OpenStudy (anonymous):

The total rainfall in Forks, Washington, this October was 8 3/8in. In September, 60% less rain fell than in October. How much rain fell in September in Forks, Washington? Express your answer in simplest form ____ _____ In. ____ @iGreen @PaulaLovesSchool13

OpenStudy (anonymous):

Hmmmm, this I don't know

OpenStudy (anonymous):

I could find a link to help though

OpenStudy (texaschic101):

8 3/8 inches -- turn to improper fraction = 67/8 or 8.375 60% less rain... 8.375 - .60(8.375) = 8.375 - 5.025 = 3.35 inches OR.. 60% less means you just paid 40% 8.375 * .40 = 3.35 inches

OpenStudy (anonymous):

but what is the last fraction i need the fraction form

OpenStudy (texaschic101):

3.35 = 3 35/100 = 3 7/20

OpenStudy (anonymous):

thanks soo much

OpenStudy (texaschic101):

no problem :)